OpenStudy (anonymous):

I will fan and give a medal if someone helps me with Pythagorean Theorem please!!!

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

give me a sec please

OpenStudy (anonymous):

ok sorry no rush

OpenStudy (anonymous):

if x =12 and y =16 what is the length of z?? A. 400 B.20 C.112 D.4

OpenStudy (anonymous):

Pythagorean theorem is:\[\Large x^2+y^2=z^2\]Plug in 12 for x and 16 for y.\[\Large (12)^2+(16^2)=z^2\]12^2 = 144 and 16^2 = 256\[\Large144+256=z^2\]\[\Large z^2=400\]Take the square root of both sides.\[\Large z=20\]So C.

OpenStudy (the_fizicx99):

The formula is a^2 + b^2 = c^2 12^2 + 16^2 = 400 c^2 = 400 c = 20

OpenStudy (anonymous):

will u help me with like 9 more pleaseeeeeee

OpenStudy (the_fizicx99):

So B.

OpenStudy (anonymous):

The city of Nairobi, Kenya, is 470 km east and 185 km south of the city of Kampala, Uganda. If Matunde is flying from Nairobi to Kampala, what is the shortest possible distance of his flight? (Round to the nearest km.) A.505 km B.432 km C.563 km D.655 km

OpenStudy (anonymous):

Oops I got 20 but I said C for some reason lol

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

470^2 + 185^2 = c^2 Same idea as previous one so I'll skip that stuff A 505 km

OpenStudy (anonymous):

if x=61 cm and y=3 cm what is the length of side z?

OpenStudy (anonymous):

@bagajr

OpenStudy (anonymous):

do u get it @tHe_FiZiCx99

OpenStudy (the_fizicx99):

61